A rectangular window pane is 12 square feet. The width of the window pane is 1 foot less than its length.
Find the width of the window pane.

Answer by amarjeeth123(571)

Let the length of the window pane be x
Width of the window pane is (x-1).
Area of the window pane=x(x-1)
x(x-1)=12
x^2-x-12=0
x^2-4x+3x-12=0
x(x-4)+3(x-4)=0
(x-4)(x+3)=0
x=4
Width of the window pane is 3 feet.